# Unified interpretation of 95 GeV di-photon and di-tau Excesses in the Georgi-Machacek Model

### v1

@@ 4 Methodology and Constraints
In order to identify viable regions of parameter space within the GM model, we perform a numerical scan subject to both theoretical and experimental constraints. This section describes the methodology of the scan together with the various constraints imposed. The detailed numerical outcomes corresponding to these constraints will be presented in Sec. 5.
4.1 Experimental constraints
As discussed in Sec. 2, the GM model extends the SM by introducing several new scalar particles. Therefore, its predictions must be consistent with current experimental observations.
  • Electroweak precision measurements.  The effects of new scalar states in the GM model enter electroweak precision observables through their contributions to the self-energies of the $W$ and $Z$ bosons as well as neutral-current processes. At the one-loop level such oblique corrections are conveniently parametrized by the Peskin-Takeuchi variables $S$, $T$, and $U$ Peskin:1990zt ; Peskin:1991sw . The latest global fit based on precision electroweak data yields ParticleDataGroup:2024cfk
    $T=0.01\pm 0.12,\qquad S=-0.04\pm 0.10,\qquad U=-0.01\pm 0.09.$ (29)
@@ 4.1 Experimental constraints
    Since custodial symmetry is preserved at tree level in the GM model due to its elegant scalar field structure, the slight deviation provides an intriguing hint about the GM model.
  • Higgs data fits.   The heavier singlet scalar $h$, identified as the SM-like Higgs boson at 125 GeV, must remain consistent with the combined ATLAS and CMS measurements at the $95\%$ confidence level. This requirement is implemented through package HiggsSignals-2.6.2 Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn ; Bahl:2022igd , demanding a $p$-value greater than $0.05$ and a total $\chi^{2}$ not exceeding the SM value by more than $\Delta\chi^{2}<6.18$ (corresponding to $95\%$ C.L. for two degrees of freedom).
    The couplings of the 125 GeV Higgs boson to fermions and vector bosons are also probed by global fits in the $\kappa$-framework. The ATLAS analysis reports
    $\begin{split}&\kappa_{tt}=0.94\pm 0.11,\quad\kappa_{bb}=0.89\pm 0.11,\quad\kappa_{\tau\tau}=0.93\pm 0.07,\\ &\kappa_{\mu\mu}=1.06^{+0.25}_{-0.30},\quad\kappa_{WW}=1.05\pm 0.06,\quad\kappa_{ZZ}=0.99\pm 0.06,\end{split}$ (31)
    while the CMS fit gives
    $\begin{split}&\kappa_{tt}=1.01\pm 0.11,\quad\kappa_{bb}=0.99\pm 0.16,\quad\kappa_{\tau\tau}=0.92\pm 0.08,\\ &\kappa_{\mu\mu}=1.12\pm 0.21,\quad\kappa_{WW}=1.02\pm 0.08,\quad\kappa_{ZZ}=1.04\pm 0.07.\end{split}$ (32)
    Combining the two sets of results, the approximate $1\sigma$ ranges adopted in the scan are
    $\begin{split}0.83\leq\kappa_{tt}\leq 1.12,\quad 0.78\leq\kappa_{bb}\leq 1.15,\quad 0.84\leq\kappa_{\tau\tau}\leq 1.00,\\ 0.81\leq\kappa_{\mu\mu}\leq 1.36,\quad 0.94\leq\kappa_{WW}\leq 1.11,\quad 0.93\leq\kappa_{ZZ}\leq 1.11.\end{split}$ (33)
    and these constraints are implemented in the scan through package HiggsSignals.
    In addition, the production and decay signal strengths of any extra Higgs states must not violate collider exclusion limits. These bounds, derived from LEP and LHC searches, are applied using HiggsBounds-5.10.0 Bahl:2022igd ; Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv , ensuring consistency with current experimental data.
  • $B$-physics observables. Rare decays of $B$ mesons provide sensitive probes of physics beyond the SM, since such processes are strongly suppressed within the SM itself. In particular, the branching ratios of $B_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$ are well measured and offer stringent constraints on extended Higgs sectors CMS:2014xfa . In our analysis we require
    $\begin{split}1.7\times 10^{-9}<&\mathcal{B}(B_{s}\to\mu^{+}\mu^{-})<4.5\times 10^{-9},\\ 2.99\times 10^{-4}<&\mathcal{B}(B\to X_{s}\gamma)<3.87\times 10^{-4},\end{split}$ (34)
    corresponding to the current $2\sigma$ experimental intervals.
4.2 Theoretical constraints
In addition to the above experimental limits, theoretical considerations further constrain the parameter space, ensuring the internal consistency of the GM model.
@@ 4.2 Theoretical constraints
  • Perturbative unitarity.  Perturbative unitarity requires that the zeroth partial-wave amplitude $a_{0}$ of scalar-scalar scattering obey $|a_{0}|<1$, or equivalently ${\rm Re}(a_{0})<1/2$. In the GM model this translates into bounds on linear combinations of quartic couplings Hartling:2014zca ; Chowdhury:2024mfu ; Aoki:2007ah , leading to the conditions
    $\begin{split}\sqrt{(6\lambda_{1}-7\lambda_{3}-11\lambda_{4})^{2}+36\lambda_{2}^{2}}+\left|6\lambda_{1}+7\lambda_{3}+11\lambda_{4}\right|&<4\pi,\\ \sqrt{(2\lambda_{1}+\lambda_{3}-2\lambda_{4})^{2}+\lambda_{5}^{2}}+\left|2\lambda_{1}-\lambda_{3}+2\lambda_{4}\right|&<4\pi,\\ \left|2\lambda_{3}+\lambda_{4}\right|<\pi,\qquad\left|\lambda_{2}-\lambda_{5}\right|&<2\pi.\end{split}$ (35)
    These inequalities are applied to all scanned points to guarantee the perturbative consistency of the scalar sector.
@@
  • Positive definiteness constraints.  It is well known that quantum corrections can significantly modify the shape of the scalar potential, potentially leading to violations of the tree-level bounded-from-below (BFB) conditions Hartling:2014zca . Such effects may induce deeper vacua at large field values or render the potential unbounded, implying that the electroweak vacuum preserving custodial symmetry may not be the true ground state. This situation is reminiscent of the metastability of the SM electroweak vacuum for a 125 GeV Higgs boson.
@@
    To address this issue, we impose the so-called positive definiteness constraints Du:2024mry , which guarantee the stability of the custodial vacuum by requiring positivity of certain subdeterminants of the quartic coupling matrices in the one-loop RG-improved potential. We parameterize the one-loop RG-improved scalar potential by the dimensionless couplings $\lambda$, $\rho_{1,2,\dotsc,5}$ and $\sigma_{1,2,3,4}$ of the most general ${\rm SU(3)}_{C}\times{\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$ invariant GM model. [^2] Dimensional parameters are eliminated via the minimisation conditions for the VEVs, and thus do not affect the large-field analysis relevant to positive definiteness. In the custodial-symmetric limit these couplings map onto the standard GM quartics $\lambda_{1,2,\dots,5}$ as
@@
    where the $\mathcal{V}$ defined as
@@
    and $\mathcal{V}_{n}$ denotes the $n\times n$ upper-left submatrix of $\mathcal{V}$.
    It should be noted that the coupling parameters (such as $\lambda$, $\rho_{1,2,\dotsc,5}$, and $\sigma_{1,2,3,4}$, etc.) must be replaced by their renormalization group evolved values in regions of large field values. These evolved parameters can then be used to determine the vacuum structure of the GM model.
    We utilise the iterative method described in Du:2024mry to assess whether the parameter space that adequately explains the observed excess signals satisfies the positive definiteness constraints. The gauge coupling values at the EW scale employed in this study are derived from Antusch:2013jca . To ensure that the vacuum maintains custodial symmetry, the boundary conditions for the scalar quartic couplings at the EW scale, which are crucial for the RGEs, are selected in accordance with Eq. (36). To determine the scale dependence of the quartic couplings within the one-loop RG-improved tree-level scalar potential, it is essential to evolve the input parameters to an arbitrary scale, especially in regions of large field values, using the appropriate beta functions. In our detailed numerical analysis, to establish the positive definiteness constraints, we evolve the relevant quartic couplings and mass dimension parameters from the EW scale ($M_{Z}=91.2~{\rm GeV}$) up to the Planck scale ($M_{\rm Planck}\approx 1.12\times{10}^{19}~{\rm GeV}$) or to the lowest Landau pole scale, where any of the quartic couplings diverge.
4.3 Scanning setup
In order to realise the light custodial singlet scalar $H$ lie near $95.4~{\rm GeV}$ capable of accomodating the LHC hints in the $\gamma\gamma$ and $\tau\tau$ channels. In practice, we ultilise the package GMCALC Hartling:2014xma , a specifical GM model calculator. Alternatively, we use the SPheno package Porod:2003um ; Porod:2011nf in combination with SARAH Staub:2013tta ; Staub:2012pb ; Staub:2010jh ; Staub:2009bi ; Staub:2008uz to compute the GM model spectrum. We use the package HiggsSignals-2.6.2 Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn ; Bahl:2022igd to examine whether the properties of the next-to-lightest Higgs boson $h$ resemble those of the SM-like Higgs boson. The additional light Higgs boson $H$ is tested against collider constraints using the package HiggsBounds-5.10.0 Bahl:2022igd ; Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv .
We conduct a comprehensive scan of the input parameters in Eq. (20), applying all constraints described in Sec. 4.1 and Sec. 4.2, over the following ranges:
$\begin{split}-\sqrt{4\pi}&<{\lambda_{2,3,4,5}}<\sqrt{4\pi},\\ 50~\mathrm{GeV}&<m_{h,H}^{\rm In}<200~\mathrm{GeV},\\ 0<\sin{\theta}&<0.45,\quad 0<\alpha<\pi.\end{split}$ (39)
The two singlet Higgs states are restricted to
$94.4~{\rm GeV}<m_{H}<96.4~{\rm GeV},\quad 124.1~{\rm GeV}<m_{h}<126.1~{\rm GeV}.$ (40)
The $\chi^{2}$ function quantifying the ability of light custodial singlet scalar $H$ to explain the excesses in the $\gamma\gamma$ and $\tau\tau$ channels is given by
$\chi^{2}_{95}=\left(\frac{\mu_{\gamma\gamma}-0.24}{0.09}\right)^{2}+\left(\frac{\mu_{\tau\tau}-1.2}{0.5}\right)^{2}.$ (41)
In the absence of a published combined likelihood, correlations between $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are neglected. Unless stated otherwise, points with $\chi^{2}_{95}<6.18$ (two observables) are regarded as consistent at the $2\sigma$ level.

### v2

@@ 4 Methodology and Constraints
In order to realize a light custodial singlet scalar $H$ near $95.4~{\rm GeV}$ that is capable of accommodating the LHC hints in the $\gamma\gamma$ and $\tau\tau$ channels, we perform a comprehensive parameter scan. This section describes the methodology of the scan together with the various constraints imposed. The detailed numerical outcomes corresponding to these constraints will be presented in Sec. 5.
In practice, we utilize the package GMCALC Hartling:2014xma , a dedicated GM model calculator. Alternatively, we use the SPheno package Porod:2003um ; Porod:2011nf in combination with SARAH Staub:2013tta ; Staub:2012pb ; Staub:2010jh ; Staub:2009bi ; Staub:2008uz to compute the GM model spectrum.
We perform a comprehensive scan of the input parameters in Eq. (20) over the following ranges:
$\begin{split}-\sqrt{4\pi}&<{\lambda_{2,3,4,5}}<\sqrt{4\pi},\\ 120~\mathrm{GeV}<m_{h}^{\rm In}&<130~\mathrm{GeV},\quad 90~\mathrm{GeV}<m_{H}^{\rm In}<102~\mathrm{GeV},\\ 0<\sin{\theta}&<0.4,\quad\pi/4<|\alpha|<\pi/2.\end{split}$ (29)
To assess the global viability of the GM model in interpreting the observed excesses, we construct a likelihood-based framework. The total likelihood $\mathcal{L}_{\rm tot}$ is defined as
$\mathcal{L}_{\rm tot}\,=\,\exp\!\left(-\frac{1}{2}\,\chi^{2}_{\rm tot}\right)\times\Theta_{\rm PD}\,\Theta_{\rm unit}\,\Theta_{\tt HB}\,\Theta_{\tt HS},\qquad\chi^{2}_{\rm tot}=\chi^{2}_{\rm mass}+\chi^{2}_{95}+\chi^{2}_{\rm EW}+\chi^{2}_{\kappa}+\chi^{2}_{B}.$ (30)
The total $\chi^{2}_{\rm tot}$ contains only genuine likelihood contributions from the experimental observables, while the factors $\Theta_{\rm PD}$, $\Theta_{\rm unit}$, $\Theta_{\tt HB}$, and $\Theta_{\tt HS}$ act as acceptance criteria that restrict the likelihood support to theoretically and phenomenologically viable regions of parameter space.
The individual $\chi^{2}$ and $\Theta$ terms entering Eq. (30) are defined as follows.
  • $\chi^{2}_{\rm mass}=\left(\dfrac{m_{h}-125.18~\rm{GeV}}{1~\rm{GeV}}\right)^{2}+\left(\dfrac{m_{H}-95.4~\rm{GeV}}{1~\rm{GeV}}\right)^{2}$.
    Here, $m_{h}$ and $m_{H}$ are the theoretical predictions of the two $CP$-even Higgs states, and the total uncertainties (theoretical and experimental) are both set to $1~{\rm GeV}$.
  • $\Theta_{\tt HB}$ implements the direct-search constraints from HiggsBounds-5.10.0 Bahl:2022igd ; Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv . We define $\Theta_{\tt HB}=1$ for parameter points that satisfy the HiggsBounds exclusion test and $\Theta_{\tt HB}=0$ otherwise. In this way, direct-search constraints act as a veto on the parameter space rather than as a compensating contribution to the global $\chi^{2}$.
  • $\Theta_{\tt HS}$ encodes the consistency requirement from HiggsSignals-2.6.2 Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn ; Bahl:2022igd . To avoid double counting Higgs signal-strength information with the explicit $\chi^{2}_{\kappa}$ term defined below, we do not include the HiggsSignals output as an additive likelihood contribution in $\chi^{2}_{\rm tot}$. Instead, we require that the sample satisfy $\chi^{2}({\rm GM})<\chi^{2}({\rm SM})$, where $\chi^{2}({\rm SM})\approx 152.5$ denotes the SM reference value returned by HiggsSignals for the 125 GeV Higgs rate measurements in our setup. We therefore set $\Theta_{\tt HS}=1$ when this condition is fulfilled and $\Theta_{\tt HS}=0$ otherwise.
  • $\chi^{2}_{95}=\left(\dfrac{\mu_{\gamma\gamma}-0.24}{0.09}\right)^{2}+\left(\dfrac{\mu_{\tau\tau}-1.2}{0.5}\right)^{2}$: denotes the two $95~{\rm GeV}$ excesses, where $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are the theoretical predictions including loop corrections.
  • The contribution $\chi^{2}_{\rm EW}$ encodes constraints from electroweak precision measurements in terms of the oblique parameters $S$, $T$, and $U$, which capture the one-loop contributions of the additional scalar states to gauge-boson self-energies and neutral-current processes Peskin:1990zt ; Peskin:1991sw . In this work, the $\chi^{2}_{\rm EW}$ is implemented using a multivariate Gaussian form,
    $\chi^{2}_{\rm EW}=\Delta\vec{O}_{\rm EW}^{\,T}\,\rho_{\rm EW}^{-1}\,\Delta\vec{O}_{\rm EW},$ (31)
    where
    $\Delta\vec{O}_{\rm EW}^{T}=\left(\frac{S+0.04}{0.10},\;\frac{T-0.01}{0.12},\;\frac{U-0.05}{0.09}\right)$ (32)
    and $\rho_{\rm EW}$ represents the correlation matrix, defined as
    $\rho_{\rm EW}=\begin{pmatrix}1&0.93&-0.70\\[4.0pt] 0.93&1&-0.87\\[4.0pt] -0.70&-0.87&1\end{pmatrix}$. (33)
@@
    In the custodial-symmetric GM setup, $\rho_{\rm tree}=1$ holds by construction, and any residual deviation arises only at loop level.
  • The contribution $\chi^{2}_{\kappa}$ encodes the constraints on the Higgs coupling modifiers through a two-dimensional Gaussian likelihood in the $(\kappa_{V},\kappa_{F})$ plane, and is defined as
    $\chi^{2}_{\kappa}=\Delta\vec{\kappa}^{\,T}\,V^{-1}\,\Delta\vec{\kappa},$ (35)
    where
    $\Delta\vec{\kappa}=\begin{pmatrix}\kappa_{V}-\hat{\kappa}_{V}\\ \kappa_{F}-\hat{\kappa}_{F}\end{pmatrix}$, (36)
    and $V$ denotes the corresponding covariance matrix,
    $V=\begin{pmatrix}\sigma_{V}^{2}&0.39\,\sigma_{V}\sigma_{F}\\ 0.39\,\sigma_{V}\sigma_{F}&\sigma_{F}^{2}\end{pmatrix}$. (37)
    In our analysis, we adopt the ATLAS combined fit results shown in Fig. 4 of Ref. ATLAS:2022vkf , with $\hat{\kappa}_{V}=1.035\pm 0.031$ and $\hat{\kappa}_{F}=0.95\pm 0.05$, and a correlation coefficient of $0.39$.
  • $\chi_{B}^{2}=\left(\dfrac{\mathcal{B}(B_{s}\to\mu^{+}\mu^{-})-3.1\times{10^{-9}}}{1.3\times{10^{-9}}}\right)^{2}+\left(\dfrac{\mathcal{B}(B\to X_{s}\gamma)-3.43\times{10^{-4}}}{0.44\times{10^{-4}}}\right)^{2}$ accounts for constraints from rare $B$-meson decays, which are strongly suppressed in the SM and therefore provide sensitive probes of physics beyond the SM, in particular extended Higgs sectors CMS:2014xfa .
In addition to the experimental limits mentioned above, theoretical considerations further constrain the parameter space through the Heaviside step functions $\Theta_{\rm unit}$ and $\Theta_{\rm PD}$ in Eq. (30). These functions act as strict vetoes, restricting the likelihood support to regions of parameter space that are theoretically consistent, where $\Theta(x)$ equals one for $x\geq 0$ and zero otherwise.
@@
The $\Theta_{\rm unit}$ enforces perturbative unitarity of scalar $2\to 2$ scattering, requiring that the zeroth partial-wave amplitudes obey $|{\rm Re}\,a_{0}|<\tfrac{1}{2}$ (equivalently $|a_{0}|<1$). In the GM model this translates into bounds on linear combinations of quartic couplings Hartling:2014zca ; Chowdhury:2024mfu ; Aoki:2007ah , leading to the conditions
$\begin{split}\sqrt{(6\lambda_{1}-7\lambda_{3}-11\lambda_{2})^{2}+36\lambda_{4}^{2}}+\left|6\lambda_{1}+7\lambda_{3}+11\lambda_{2}\right|&<4\pi,\\ \sqrt{(2\lambda_{1}+\lambda_{3}-2\lambda_{2})^{2}+\lambda_{5}^{2}}+\left|2\lambda_{1}-\lambda_{3}+2\lambda_{2}\right|&<4\pi,\\ \left|2\lambda_{3}+\lambda_{2}\right|<\pi,\qquad\left|\lambda_{4}+\lambda_{5}\right|&<2\pi.\end{split}$ (38)
Therefore, $\Theta_{\rm unit}=1$ if all conditions in Eq. (38) are satisfied simultaneously.
@@
The positive-definiteness requirement is imposed to ensure the stability of the electroweak vacuum, such that the scalar potential remains bounded from below for large field values. The quantum corrections can significantly modify the shape of the scalar potential, potentially leading to violations of the tree-level bounded-from-below (BFB) conditions Hartling:2014zca . Such effects may induce deeper vacua at large field values or render the potential unbounded, implying that the electroweak vacuum preserving custodial symmetry may not be the true ground state. This situation is reminiscent of the metastability of the SM electroweak vacuum for a $125~{\rm GeV}$ Higgs boson.
@@
To address this issue, we impose the positive definiteness constraints Du:2024mry to guarantee the stability of the custodial vacuum by requiring positivity of certain subdeterminants of the quartic coupling matrices in the one-loop RG-improved potential. For simplicity, the one-loop RG-improved scalar potential is parameterized by the dimensionless couplings $\lambda$, $\rho_{1,2,\dotsc,5}$, and $\sigma_{1,2,3,4}$ of the most general ${\rm SU(3)}_{C}\times{\rm SU(2)}_{L}\times{\rm U(1)}_{Y}$ invariant GM model. [^2] Dimensional parameters are eliminated via the minimization conditions for the VEVs, and thus do not affect the large-field analysis relevant to positive definiteness. In the custodial-symmetric limit these couplings map onto the standard GM quartics $\lambda_{1,2,\dots,5}$ as
@@
where $\mathcal{V}$ is defined as
@@
and $\mathcal{V}_{n}$ denotes the $n\times n$ upper-left submatrix of $\mathcal{V}$. Therefore, the $\Theta_{\rm PD}$ in Eq. (30) is defined as
$\Theta_{\rm PD}=\prod_{n}\Theta(\det\mathcal{V}_{n}).$ (42)
In practice, the full matrix is used numerically, while the analytic form shown here corresponds to the simplified limit.
It should be noted that the coupling parameters (such as $\lambda$, $\rho_{1,2,\dotsc,5}$, and $\sigma_{1,2,3,4}$, etc.) must be replaced by their renormalization group evolved values in regions of large field values. These evolved parameters can then be used to determine the vacuum structure of the GM model. We utilize the iterative method described in Du:2024mry to assess whether the parameter space that adequately explains the observed excess signals satisfies the positive definiteness constraints. The gauge coupling values at the EW scale employed in this study are derived from Antusch:2013jca . To ensure that the vacuum maintains custodial symmetry, the boundary conditions for the scalar quartic couplings at the EW scale, which are crucial for the RGEs, are selected in accordance with Eq. (39). To determine the scale dependence of the quartic couplings within the one-loop RG-improved tree-level scalar potential, it is essential to evolve the input parameters to an arbitrary scale, especially in regions of large field values, using the appropriate beta functions. In our detailed numerical analysis, to establish the positive definiteness constraints, we evolve the relevant quartic couplings and mass dimension parameters from the EW scale ($M_{Z}=91.2~{\rm GeV}$) up to the Planck scale ($M_{\rm Planck}\approx 1.12\times{10}^{19}~{\rm GeV}$) or to the lowest Landau pole scale, where any of the quartic couplings diverge.
To ensure a numerically reliable assessment of the GM model’s ability to simultaneously explain the two observed excesses, we conducted an extensive exploration of the parameter space using a combination of random scans, Markov Chain Monte Carlo (MCMC) techniques, and MultiNest sampling. These complementary methods provide both broad global coverage and efficient identification of phenomenologically relevant regions with higher likelihood. In total, we sampled approximately $\mathcal{O}(10^{8})$ points, providing sufficient statistical coverage to ensure that our conclusions are robust and unaffected by scanning artifacts. Convergence was confirmed by verifying the stability of the profile likelihood (PL) with additional sampling.